OpenStudy (anonymous):

Given points P(7, 5), Q(8, 3), R(0, -1), and S(-1, 1), which statement is true? a)PQ is parallel to RS b)Pq is perpendicular to RS c) PR is perpendicular to QS d)Pr is parallel to QS

OpenStudy (dumbcow):

A use slope formula to verify

OpenStudy (anonymous):

whats the slope formula?

OpenStudy (dumbcow):

m = y2-y1/x2-x1

OpenStudy (dumbcow):

PQ = (5-3)/(7-8) = -2 RS = (-1-1)/(0-(-1)) = -2 same slope means parallel

OpenStudy (anonymous):

thanks darling (: